Solve by addition method
3/7x + 5/9y = 27
1/9x + 2/7y =7

Answer by stanbon(75887)

Solve by addition method
3/7x + 5/9y = 27
1/9x + 2/7y =7
--------------
The lcm of 7 and 9 is 63.
Multiply thru both equations by 63 to get:
------------------------
27x + 35y = 63*27
7x + 18y = 63*7
--------------------------
Multiply thru the 1st equation by 7 and the 2nd equatio by -27 to get:
7*27x + 7*35y = 7*63*27
7*-27x+18*-27y = -27*63*7
------------------------------------
Add to solve for "y":
(7*35-18*27)y = 7*63*27 - 27*63*7
y = 0
------------------
Substitute into 7x + 18y = 7*63 to solve for "x":
7x = 7*63
x = 63
=========================================
Solution: x= 63 and y = 0
=========================================
Cheers,
Stan H.
=========================================